OpenStudy (anonymous):

Can someone check my answer?? i think its B. Which pair of expressions has the same value? A. (2)(-8) and 2 -18 B. (-2)(8) and 18 + (-2) C. 2(8) and 2 + 18 D.(-2)(-8)and 2 + (-18)

OpenStudy (atlas):

No its not B (-2)8 = -2 * 8 = -16 18 +(-2) =16

OpenStudy (adrynicoleb):

Hmm. You might want to check your work.

OpenStudy (adrynicoleb):

^^

OpenStudy (anonymous):

ok thanks guys

OpenStudy (adrynicoleb):

Did you get the answer already? o.O

OpenStudy (anonymous):

D. ?

OpenStudy (anonymous):

i got D. from what i worked out

OpenStudy (adrynicoleb):

Well a negative times a negative equals a positive. So (-2)(-8) would be 16. And for 2 + (-18), the answer would be -16 because the negative is bigger than a positive.

OpenStudy (anonymous):

my last choice is C. but i dont think it can be that because i dont believe it was a match when i did it

OpenStudy (adrynicoleb):

No, it is not a match. 2(8) is 16, and 2 + 18 equals 20.

OpenStudy (adrynicoleb):

There is one more choice though *WINK WINK*

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

A. !

OpenStudy (adrynicoleb):

Lol yes. Can you show me why it's A?

OpenStudy (anonymous):

-2 times 8 =-16 2 - 18=-16

OpenStudy (adrynicoleb):

Yay!! Awesome job :)

OpenStudy (anonymous):

can u check more?

OpenStudy (anonymous):

like 3??

OpenStudy (adrynicoleb):

Are they just like this?

OpenStudy (anonymous):

no

OpenStudy (adrynicoleb):

I can try. I'm not the best at math. But open a new post so others can try to help you if I can't. :P